OpenStudy (anonymous):

Question Part Points Submissions Used Solve the following equation, giving the exact solutions which lie in [0, 2π). (Enter your answers as a comma-separated list.) tan2(x) = 3 2 sec(x)

OpenStudy (anonymous):

\[\tan^2(x)= (3/2)\sec(x)\] here is the equation better

jimthompson5910 (jim_thompson5910):

\[\large \tan^2(x)= \frac{3}{2}\sec(x)\] \[\large \sec^2(x)-1= \frac{3}{2}\sec(x)\] \[\large z^2-1= \frac{3}{2}z\] \[\large 2(z^2-1)= 3z\] \[\large 2z^2-2= 3z\] \[\large 2z^2-2-3z = 0\] \[\large 2z^2-3z-2 = 0\] Now use the quadratic formula to solve for z. Tell me what you get.

OpenStudy (anonymous):

ok sec

OpenStudy (anonymous):

I get z = -(1/2), 2

jimthompson5910 (jim_thompson5910):

so this means that sec(x) = -1/2 or sec(x) = 2 because z = sec(x)

jimthompson5910 (jim_thompson5910):

solve each equation for x

jimthompson5910 (jim_thompson5910):

one of those equations has no solutions at all

OpenStudy (anonymous):

I'm confused on the next step

jimthompson5910 (jim_thompson5910):

sec(x) = -1/2 1/cos(x) = -1/2 cos(x) = -2 what's next?

OpenStudy (anonymous):

1/cos(2)

OpenStudy (anonymous):

im guessing im wrong

jimthompson5910 (jim_thompson5910):

use the unit circle and tell me when cos(x) = -2 is true

OpenStudy (anonymous):

it doesn't exist right?

jimthompson5910 (jim_thompson5910):

good, there are no solutions

jimthompson5910 (jim_thompson5910):

sec(x) = 2 1/cos(x) = 2 cos(x) = 1/2 what are the solutions here?

OpenStudy (anonymous):

pi/3 , 5pi/3 ?

jimthompson5910 (jim_thompson5910):

good, those are your two solutions

OpenStudy (anonymous):

yay... thank you soo much!

jimthompson5910 (jim_thompson5910):

you're welcome